# VeriLLM: A Lightweight Framework for Publicly Verifiable Decentralized Inference

### v3

@@ 6.4.1 Quantization Attack
Thresholds are calibrated from benign full-precision runs. Honest executions pass $T_{\mathrm{off}}$ with probability $1-\operatorname{\textsf{negl}}(\lambda)$, whereas 8-bit quantization causes detectable statistical shifts. (ii) Lightweight on-chain spot check. The contract performs a relaxed test $T_{\mathrm{on}}$ on VRF-sampled final-token scalars, checking $|v_{j}-\mathbf{V}^{(T)}_{j}|\!\leq\!\varepsilon$ and aggregating agreement against a threshold $\theta$ [chen2024zkml]. This sustains liveness while bounding on-chain cost to $O(n)$ sampled entries.
@@
6.4.2 Early-Termination Attack
@@
Defenses. (i) Tail-token consistency. The last-segment verifier recomputes logits from the reported final hidden state and verifies that the decoding policy yields the claimed token [zhang2020delphi]. Any premature termination produces a mismatch at the tail step with probability 1 for deterministic decoding and with high probability for deterministically reseeded stochastic decoding. (ii) Per-step binding. Each segment’s boundary state is individually committed and signed. Modifying terminal states or inserting fake tokens requires forging a signature or producing an inconsistent Merkle opening, both computationally infeasible [merkle1989certified, micali1999verifiable]. Scheduler relay signatures bind token order and prevent state splicing.
@@
6.4.3 Forged Output via Small Model
@@
6.4.4 Lazy Verifier (Free-Riding)
@@
Defenses. (i) Commit–then–open discipline. Before seeing sampling indices, each verifier must commit to its final-token tensor root and later open VRF-selected positions with inclusion proofs [micali1999verifiable]. Without genuine computation, the probability of producing valid openings is negligible. (ii) Task-type indistinguishability. Because inference and verification share identical RPC interfaces and data formats, workers cannot distinguish verification probes from live inference [zhang2020delphi], preventing selective laziness. (iii) Reward gating. Rewards are paid only to verifiers whose openings pass on-chain checks; mismatches or failures trigger stake slashing. Hence, skipping computation is a strictly dominated strategy.
@@
6.5 Collusion, Censorship, and Dispute
Collusion among verifiers. A coalition may attempt to fabricate mutually consistent hidden states and corresponding Merkle roots such that on-chain openings match internally forged commitments. The commit–then–sample discipline prevents any post hoc adaptation: once a root is posted, the coalition must open VRF-chosen indices consistent with that fixed root. To ensure semantic correctness with respect to the prescribed model, VeriLLM introduces two safeguards: (i) token–state cross-checks at the tail segment, where verifiers decode from recomputed last-layer states and compare tokens to the published outputs; and (ii) a zero-knowledge escalation path, in which any single honest verifier can submit a succinct proof that a challenged prefix of the transcript violates the model’s forward mapping or decoding rule [chen2024zkml, sun2024zkllm, qu2025zkgpt]. The contract verifies the proof on-chain, finalizing the inconsistency independently of the coalition’s reports.
Censorship resistance. All commitments are recorded on-chain, and verifier selection derives from verifiable randomness. To suppress detection, an adversary would need to corrupt or block all selected verifiers for a task. Under the one-honest-verifier assumption, at least one remains live to perform offline recomputation, pass the sampled checks, or post a zero-knowledge inconsistency proof. A total denial of service is observable through missed deadlines, which activate timeout-based remediation.
6.6 Completeness and Soundness Summary
@@
$\Pr[\mathcal{E}]\leq\operatorname{\textsf{negl}}(\lambda),$ (43)
@@
7 System Implementation Details
@@
Signature and serialization pipeline. Nodes employ Ed25519 signatures for commitments and message authentication. On an RTX 5090 host, signing latency is 0.19 ms per 32 B message; verification 0.21 ms; Merkle-tree construction 1.7 µs per scalar. Table II summarizes the microbenchmarks, showing total cryptographic overhead ¡ 0.8 % of end-to-end inference time.
TABLE II: Cryptographic and Serialization Overheads in VeriLLM.
@@
8 Experiment and Evaluation
8.1 Main Results
We empirically measured the divergence of hidden states produced during inference across heterogeneous compute devices (tested on Qwen2.5-7B-Instruct [qwen2_5_technical_report]; additional models are forthcoming). Although floating-point non-determinism across GPU vendors introduces small numeric drift [gupta2015deep, wang2021characterizing], these deviations are limited in magnitude and statistically distinguishable. Consequently, the hidden-state comparison test can, with high probability, determine whether two traces originate from the same underlying model computation.
We further evaluated replacing full-precision models with quantized variants. Hidden states computed by quantized models exhibit substantially larger—and statistically significant—bias relative to the full-precision baseline, consistent with prior analyses of quantization-induced drift [banner2018post, sun2024zkllm]. Hence, simple statistical tests cleanly separate hidden states from normal inference versus those produced under precision degradation or adversarial manipulation.
TABLE III: Comparison of Hidden-State Metrics for Qwen2.5-7B-Instruct Inference on M4 and RTX 5090.
# Exact Exp Match ($|\Delta|$) Exp Mismatch ($|\Delta|$) Mean $\epsilon$
>0.2>0.2 <0.2<0.2 >5>5 <5<5
64 4 13 48 2 1 0.009
256 23 46 201 7 2 0.006
1024 87 191 802 23 7 0.005
3584 311 669 2801 89 25 0.003-0.003
TABLE IV: Comparison of Hidden-State Metrics for Qwen2.5-7B-Instruct-AWQ inference on RTX 5090 with verification on Mac M4 CPU using Qwen2.5-7B-Instruct.
# Exact Exp Match ($|\Delta|$) Exp Mismatch ($|\Delta|$) Mean $\epsilon$
>0.2>0.2 <0.2<0.2 >5>5 <5<5
64 0 39 13 7 5 -0.022
256 2 161 53 23 19 0.021
1024 7 653 201 99 71 -0.019
3584 29 2276 712 346 250 $0.014$
Bit-aware statistical comparison. Building on these results, we propose a feature-based comparison of hidden states. Given reference $S$ and candidate $\bar{S}$, each floating-point element is decomposed into sign, exponent, and mantissa [wang2021characterizing]. With tolerances $e_{w}>0$ (e.g., $e_{w}=10^{-2}$) and $e_{m}\!\in\!\mathbb{N}$ (e.g., 7 mantissa steps), we define:
1. Exponent-first checks. $P_{e}$ — proportion of exponent mismatches; $P_{m}$ — fraction of large mantissa deviations among exponent mismatches; $P_{w}$ — fraction of small mantissa deviations among exponent matches. 2. Mean discrepancy. $e=\frac{1}{|\Delta|}\sum_{\delta\in\Delta}\delta$
Empirically calibrated acceptance thresholds are:
Off-chain: $P_{e}<0.05,\,P_{m}>0.75,\,P_{w}>0.80,\,e\in[-0.01,0.01]$. On-chain: $P_{e}<0.08,\,P_{m}>0.70,\,P_{w}>0.75,\,e\in[-0.02,0.02]$.
These rules balance detection sensitivity with tolerance for benign noise and are consistent with prior verifiable ML calibration studies [zhang2020delphi, chen2024zkml].
8.2 Performance and Overhead Breakdown
We evaluate VeriLLM across five dimensions: throughput and latency, on-chain cost, scalability, ablation, and failure injection [bonneau2015sok]. All tests use the 10-node setup described in Section VI.
8.2.1 Throughput and Latency
For a 7B-parameter model (512-token prompt, 128-token output), total latency per request is 1.37 s: inference 93.1 %, scheduler relay 3.4 %, and verification 3.5 %. Verification overhead averages 0.78 % of total inference time, consistent with lightweight cryptographic pipelines [micali1999verifiable]. Increasing verifier parallelism beyond three yields diminishing returns as GPU prefill already saturates compute throughput.
TABLE V: Latency Breakdown for Qwen2.5-7B-Instruct (128-token output).
Component Mean (ms) Share (%) Cumulative (%)
Prefill + Decode (Inference) 1278 93.1 93.1
Scheduler Relay / Signatures 46 3.4 96.5
Verifier Prefill (Parallel) 33 2.4 98.9
On-chain Commit + Reveal 15 1.1 100.0
8.2.2 On-chain Cost Analysis
Each inference produces two transactions: (i) commit (108 k gas, 2.9 kB) and (ii) reveal (134 k gas, 3.2 kB). Merkle proofs require 8–10 hashes per sample, yielding $\approx$ 0.004 USD per verification at current Ethereum L2 prices [chainlinkvrf2023, ethgas2024]. VRF verification adds ¡ 4 % of total cost.
8.2.3 Scalability Experiments
Verification cost per token decreases inversely with verifier count $m$ until m>6m\!>\!6, where on-chain aggregation dominates. Prefill parallelization across $L$ segments achieves near-linear scaling $T_{\text{verify}}\!\approx\!T_{\text{prefill}}/L$, consistent with prior verifiable-inference frameworks [arun2025verde].
8.2.4 Ablation Study
8.2.5 Failure Injection and Detection Latency
We emulate four attack classes—quantization, early termination, forged output, and lazy verification—via runtime perturbations [zhang2020delphi, sun2024zkllm]. Table VI summarizes detection latency (from deviation to slash). All attacks are detected and finalized within 1–3 s, confirming that the commit–sample–reveal protocol yields rapid, tunable defense.
TABLE VI: Detection latency under injected attacks.
Attack Type Detection Probability Mean Detection Latency (s)
Quantization (W8A8) $99.9\%$ 1.7
Early Termination $100\%$ 1.2
Forged Output $99.8\%$ 2.4
Lazy Verification $99.9\%$ 2.8
9 Discussion and Limitations
10 Conclusion
This work presented VeriLLM, a publicly verifiable framework for decentralized large language model (LLM) inference. Our system achieves correctness and accountability under the minimal one-honest-verifier assumption while maintaining verification overhead at roughly 1% of the inference cost. Through a hybrid design that combines Merkle-based commitments, VRF-driven sampling, and on-chain adjudication, VeriLLM enables any participant to audit computations without relying on a trusted majority or centralized coordinator.
We formalized the verification protocol, analyzed its security properties, and proved that malicious deviations—such as quantization, early termination, or forged outputs—are detected with overwhelming probability. Our implementation demonstrates that cross-device floating-point noise remains well bounded and statistically separable, allowing lightweight hidden-state comparison to distinguish honest inference from tampered or degraded computations. Experimental evaluation confirms that both off-chain and on-chain checks operate efficiently, supporting scalable deployment in heterogeneous environments.
By decoupling verification from replication and by embedding transparency into each inference trace, VeriLLM advances the practical realization of trustworthy, decentralized AI services. Future work includes integrating lightweight zero-knowledge proof systems to further compress verification proofs, extending compatibility to multimodal models, and exploring decentralized scheduling and incentive mechanisms to enhance robustness and fairness across larger networks.
Acknowledgments
The authors would like to express their sincere gratitude to the Gradient Network for its generous support. The resources, technical assistance, and constructive feedback provided by Gradient were instrumental in the successful completion of this research.
A Game-Theoretic Proof
.1 Modeling and Assumption
@@ A Game-Theoretic Proof
The protocol requires the $n$ verifiers to run the same inference task $T$ for verification, and honest verification incurs a computational cost of $c$. The inference result lies in a metric space $\Omega$ with a distance function $d:\Omega^{2}\to[0,+\infty)$ that satisfies symmetry and triangle inequality. While the task has an hidden ground-truth result $x=x(T)$, the inference process has a computational noise so that the inference result of each verifier has a small random noise.
@@
We assume that there exists a parameter $\delta>0$, such that for each verifier $i$, the result of honest verification is a random variable $x_{i}$ s.t. $P(d(x_{i},x)<\delta)>1-\epsilon_{1}$, i.e., the result lies within a small ball centered at $x$ with high probability. On the other hand, if verifier $i$ does not perform the honest verification, we assume that he can only do some random guess and report a $\tilde{x}_{i}$, and $P(d(\tilde{x}_{i},x)<3\delta)<\epsilon_{2}$, i.e., he only has a small probability to guess a result within a 3-time radius.
.2 Algorithm
@@
We let every verifier $i\in[n]$ independently report their verification results as $x_{i}$. Then we run the Kruskal algorithm until the edge length exceeds $2\delta$, getting a forest with vertex set $[n]$. For each connected component $C\subseteq[n]$, we call $C$ as a cluster in agreement.
@@
For a given quorum-size $q>\frac{n}{2}$, we check if there is a cluster in agreement of size at least $q$, called as a “proper cluster”. Because $q>\frac{n}{2}$, there can be at most one such proper cluster.
@@
If there is one, the system reaches a consensus that this cluster is accepted, and verifiers outside the cluster are rejected. We accept the inferencer if and only if his result is within $2\delta$ distance of a verifier in the proper cluster.
@@
If there is no proper cluster, then the system fails to reach a consensus and invokes the zero-knowledge proof stage. If the inferencer or any verifier disputes about the results, the zero-knowledge proof is also invoked.
.3 Analysis
In the analysis, we let $n=6,q=4,\epsilon_{1}=\epsilon_{2}=0.01,r=0.8$. In the calculation we only consider the probability that “the consensus is reached in the voting (i.e., there is a proper cluster) and the verifier is accepted”. As we assume that if the consensus is not reached and the zero-knowledge proof is invoked, the system will get the correct judgment, the actual result will be better even without the disputing process.
.3.1 Acceptance Rate of Honest Verifiers
We assume that the verifier $1$ is honest, then $P(d(x_{1},x)<\delta)>1-\epsilon_{1}$.
For verifiers $2,3,\cdots,n$, each one has a probability $r$ to be honest, and each honest verifier has a probability at least $1-\epsilon_{1}$ to report a result in the ball $B(x,\delta)$. Hence, each of them has a probability at least $p_{in}=(1-\epsilon_{1})r$ to report a result in $B(x,\delta)$.
We can see that if verifier $1$ and at least q1q-1 other verifiers’ reports lie in $B(x,\delta)$, then they are within $2\delta$ distance of each other, hence verifier $1$ belongs to a proper cluster and is accepted. Now we estimate the probability of this event.
For the n1n-1 other verifiers, the probability that at least q1q-1 lie in $B(x,\delta)$ is at least $\sum_{k=q-1}^{n-1}\binom{n-1}{k}p_{in}^{k}(1-p_{in})^{n-1-k}$. Hence, the probability that verifier $1$ is accepted is at least
$(1-\epsilon_{1})\cdot\sum_{k=q-1}^{n-1}\binom{n-1}{k}p_{in}^{k}(1-p_{in})^{n-1-k}.$
When $n=6,q=4,\epsilon_{1}=0.01,r=0.8$, we compute that the probability is greater than $92.6\%$.
.3.2 Acceptance Rate of Dishonest Verifiers
Wlog we assume that the verifier $1$ is dishonest, then $P(d(x_{1},x)<3\delta)<1-\epsilon_{2}$.
For verifiers $i\in\{2,3,\cdots,n\}$, we call $i$ regular if he reports a result in $B(x,\delta)$, and have already shown that each of them has at least an independent $p_{in}=(1-\epsilon_{1})r$ probability to be regular.
We notice that if there is no vertex in $B(x,3\delta)-B(x,\delta)$, then a cluster can only be either completely in $B(x,\delta)$, or completely outside $B(x,3\delta)$. Hence, if the dishonest verifier $1$ is accepted, at least one of following events happen:
  1. 1.
    Verifier $1$ is in a proper cluster in $B(x,\delta)$.
  2. 2.
    Verifier $1$ is in a proper cluster outside $B(x,3\delta)$.
  3. 3.
    There exists at least one vertex in $B(x,3\delta)-B(x,\delta)$.
For the first case, we know that $d(x_{1},x)<\delta<3\delta$, and the probability is at most $p_{d1}=\epsilon_{2}$.
For the second case, we know that there exist no more than nqn-q regular verifiers among $\{2,3,\cdots,n\}$. The probability is upper bounded by
$p_{d_{2}}=\sum_{k=0}^{n-q}\binom{n-1}{k}p_{in}^{k}(1-p_{in})^{n-1-k}.$
For the third case, the probability that $v_{1}$ lies in $B(x,3\delta)-B(x,\delta)$ is at most $\epsilon_{2}$, and the probability that each $v_{i}(i\geq 2)$ lies in $B(x,3\delta)-B(x,\delta)$ is at most $\epsilon_{1}r+\epsilon_{2}(1-r)$. So the total probability is upper bounded by
$p_{d_{3}}=\epsilon_{2}+(n-1)\left(\epsilon_{1}r+\epsilon_{2}(1-r)\right).$
When $n=6,q=4,\epsilon_{1}=\epsilon_{2}=0.01,r=0.8$, we can compute that $p_{d1}=0.01,p_{d2}<0.065,p_{d3}=0.06$, so the probability that a dishonest verifier gets accepted is less than $12.5\%$.

### v4

@@ A.1. Quantization Attack
Thresholds are calibrated from benign full-precision runs. Honest executions pass $T_{\mathrm{off}}$ with probability $1-\operatorname{\textsf{negl}}(\lambda)$, whereas 8-bit quantization causes detectable statistical shifts. (ii) Lightweight on-chain spot check. The contract performs a relaxed test $T_{\mathrm{on}}$ on VRF-sampled final-token scalars, checking $|v_{j}-\mathbf{V}^{(T)}_{j}|\!\leq\!\varepsilon$ and aggregating agreement against a threshold $\theta$ (Chen et al., 2024). This sustains liveness while bounding on-chain cost to $O(n)$ sampled entries.
@@
A.2. Early-Termination Attack
@@
Defenses. (i) Tail-token consistency. The last-segment verifier recomputes logits from the reported final hidden state and verifies that the decoding policy yields the claimed token (Zhang et al., 2020). Any premature termination produces a mismatch at the tail step with probability 1 for deterministic decoding and with high probability for deterministically reseeded stochastic decoding. (ii) Per-step binding. Each segment’s boundary state is individually committed and signed. Modifying terminal states or inserting fake tokens requires forging a signature or producing an inconsistent Merkle opening, both computationally infeasible (Merkle, 1989; Micali et al., 1999). Scheduler relay signatures bind token order and prevent state splicing.
@@
A.3. Forged Output via Small Model
@@
A.4. Lazy Verifier (Free-Riding)
@@
Defenses. (i) Commit–then–open discipline. Before seeing sampling indices, each verifier must commit to its final-token tensor root and later open VRF-selected positions with inclusion proofs (Micali et al., 1999). Without genuine computation, the probability of producing valid openings is negligible. (ii) Task-type indistinguishability. Because inference and verification share identical RPC interfaces and data formats, workers cannot distinguish verification probes from live inference (Zhang et al., 2020), preventing selective laziness. (iii) Reward gating. Rewards are paid only to verifiers whose openings pass on-chain consistency checks; mismatches or failures trigger stake slashing (Zhao et al., 2024). Hence, skipping computation is a strictly dominated strategy.
@@
Appendix B Discussion and Limitations
Appendix C Completeness and Soundness Summary
@@
(32) $\Pr[\mathcal{E}]\leq\operatorname{\textsf{negl}}(\lambda),$
@@
Appendix D System Implementation Details
@@
Signature and serialization pipeline. Nodes employ Ed25519 signatures for commitments and message authentication. On an RTX 5090 host, signing latency is 0.19 ms per 32 B message; verification 0.21 ms; Merkle-tree construction 1.7 µs per scalar. Table 6 summarizes the microbenchmarks, showing total cryptographic overhead ¡ 0.8 % of end-to-end inference time.
Table 6. Cryptographic and Serialization Overheads in VeriLLM.
@@
Appendix E A Game-Theoretic Proof
E.1. Modeling and Assumption
@@
The protocol requires the $m$ verifiers to run the same inference task $T$ for verification, and honest verification incurs a computational cost of $c$. The inference result lies in a metric space $\Omega$ with a distance function $d:\Omega^{2}\to[0,+\infty)$ that satisfies symmetry and triangle inequality. While the task has an hidden ground-truth result $x=x(T)$, the inference process has a computational noise so that the inference result of each verifier has a small random noise.
@@
We assume that there exists a parameter $\rho>0$, such that for each verifier $i$, the result of honest verification is a random variable $x_{i}$ s.t. $P(d(x_{i},x)<\rho)>1-\mu$, in which $\mu$ is negligibly small, i.e., the result lies within a small ball centered at $x$ with high probability. This assumption models the correctness property in the verification process: up to small random noises, the honest inference process should output consistent results.
On the other hand, if verifier $i$ does not perform the honest verification, we assume that he can only do some random guess and report a $\tilde{x}_{i}$, and $P(d(\tilde{x}_{i},x)<3\rho)<\mu$, i.e., he only has a small probability to guess a result within a 3-time radius. Here, we assume that $\mu$ is negligibly small. This assumption models the soundness property in the verification process: since all attacks in Section A will produce significantly deviated results with overwhelming probability, this criterion will capture all attacks in Section A.
E.2. Algorithm
@@
We let every verifier $i\in[m]$ independently report their verification results as $x_{i}$. Then we run the Kruskal algorithm until the edge length exceeds $2\rho$, getting a forest with vertex set $[m]$, in which $m$ is an odd number. For each connected component $C\subseteq[m]$, we call $C$ as a cluster in agreement.
@@
For a given quorum-size $q=\lceil\frac{m}{2}\rceil$, we check if there is a cluster in agreement of size at least $q$, called as a “proper cluster”. Because $q>\frac{m}{2}$, there can be at most one such proper cluster.
@@
If there is one, the system reaches a consensus that this cluster is accepted, and verifiers outside the cluster are rejected. We accept the inferencer if and only if his result is within $2\rho$ distance of a verifier in the proper cluster.
@@
If there is no proper cluster, then the system fails to reach a consensus and invokes the reverification stage with $m^{\prime}\geq m$ verifiers. If the inferencer or any verifier disputes about the results, the reverification proof is also invoked.
E.3. Analysis
Here, we first prove a lemma:
Lemma E.1.
If $r>\frac{1}{2(1-\mu)^{2}}$, and $m$ is odd, then an honest inference passes the initial verification with probability $p_{1}>\frac{1}{2}$, and a dishonest inference passes verification with probability $p_{2}<\frac{1}{2}$.
Proof.
For an honest inference $x_{0}$, from the modeling we know that $P(d(x_{0},x)<\rho)>1-\mu$, and for any honest verifier $i$, it also holds that $P(d(x_{i},x)<\rho)>1-\mu$.
Denote $\mu=r(1-\mu)$, then $\mu>\frac{1}{2(1-\mu)}>\frac{1}{2}$. Because each verifier has an independent $r$ probability to be honest, we have
(33) $p_{1}$ $=P\left(\sum_{i=1}^{m}\left[d(x_{i},x_{0})<2\rho\right]>\frac{m}{2}\right)$
(34) $\geq P\left(\sum_{i=1}^{m}\left[d(x_{i},x)<\rho\right]\cdot\left[d(x_{0},x)<\rho\right]>\frac{m}{2}\right)$
(35) $=P\left(d(x_{0},x)<\rho\right)\cdot P\left(\sum_{i=1}^{m}\left[d(x_{i},x)<\rho\right]>\frac{m}{2}\right)$
(36) $>(1-\mu)\cdot\sum_{t=\lceil\frac{m}{2}\rceil}^{m}\binom{m}{t}(r(1-\mu))^{t}(1-r(1-\mu))^{m-t}$
(37) $=(1-\mu)\cdot\sum_{t=\lceil\frac{m}{2}\rceil}^{m}\binom{m}{t}\mu^{t}(1-\mu)^{m-t}.$
From the Condorcet’s jury theorem, we have
(38) $\mu>\frac{1}{2}\implies\sum_{t=\lceil\frac{m}{2}\rceil}^{m}\binom{m}{t}\mu^{t}(1-\mu)^{m-t}\geq\mu.$
Hence, it holds that
(39) $p_{1}$ $>(1-\mu)\cdot\mu$
(40) $>(1-\mu)\cdot\frac{1}{2(1-\mu)}$
(41) $=\frac{1}{2}.$
We can also show $p_{2}<\frac{1}{2}$ by symmetry.
Then, we can find an $m^{\prime}$ such that the re-verification returns a correct result with overwhelming probability, via the following lemma:
Lemma E.2.
For any given $\xi\in(0,\frac{1}{2})$, if $m^{\prime}>\frac{\ln\frac{1}{\xi}}{2\left(r(1-\mu)-\frac{1}{2}\right)^{2}}$, then the re-verification returns a correct result with probability at least $1-\xi$.
Proof.
From Hoeffding’s equality, we have
(42) $\sum_{t=\lceil\frac{m^{\prime}}{2}\rceil}^{m^{\prime}}\binom{m^{\prime}}{t}\mu^{t}(1-\mu)^{m^{\prime}-t}$ $\geq 1-e^{-2m^{\prime}(\mu-\frac{1}{2})^{2}}$
(43) $=1-e^{\ln\xi}$
(44) $=1-\xi.$
With the two lemmas above, we can design an incentive rule in which, a dispute incurs a cost to the inferencer and a successful dispute rewards the inferencer $2$ times the cost (compared to not disputing), so that an honest inferencer is incentivized to dispute when getting rejected, but a dishonest inferencer is never incentivized to dispute. An example incentive rule is:
  • A rejected inferencer is eligible to dispute at most once. The first verification samples $m$ verifiers, and the second verification samples $m^{\prime}$ verifiers according to Lemma E.2.
  • If a inference is accepted, the inferencer and all verifiers in the proper cluster are rewarded $6$, and all disagreeing verifiers are slashed $2$.
  • If a inference is rejected and is not disputed, all verifiers in the proper cluster are rewarded $6$, and the inferencer and all disagreeing verifiers are slashed $2$.
  • To dispute incurs a cost of $8$ for the inferencer. If the dispute is successful, the cost is refunded (finally rewarded $6$); otherwise, the inferencer pays the cost and the slash (rewarded 10-10 in total). Verifiers in both rounds are rewarded $6$ or 2-2 similarly as above, according to whether they agree with the proper cluster.
Appendix F Threshold Sensitivity Analysis
To validate our threshold selection, we conduct a sensitivity analysis across all four verification metrics using results from honest inference (homogeneous and cross-platform FP32), quantization attacks, and lazy verifier scenarios. For each parameter, we vary its value while holding others fixed at their selected values.
Table 7 presents the results. The selected thresholds ($P_{e}<0.05$, $|\bar{\varepsilon}|<0.01$, $P_{m}>0.75$, $P_{w}>0.80$) achieve 99.85% attack detection with only 1.9% false positive rate. Stricter thresholds increase FPR, while looser thresholds reduce detection accuracy.
Table 7. Threshold Sensitivity Analysis. The selected thresholds achieve 99.8% attack detection rate with only 1.9% false positive rate.
Threshold FPR FNR TPR TNR
(a) $P_{e}$ Sensitivity ($|\varepsilon|<0.01$, $P_{m}>0.75$, $P_{w}>0.80$ fixed)
   $P_{e}<0.02$ 5.2% 0.03% 99.97% 94.8%
   $P_{e}<0.03$ 2.3% 0.08% 99.92% 97.7%
   $P_{e}<0.05$ (selected) 1.9% 0.15% 99.85% 98.1%
   $P_{e}<0.08$ 1.5% 0.28% 99.72% 98.5%
   $P_{e}<0.10$ 1.4% 0.41% 99.59% 98.6%
(b) $|\varepsilon|$ Sensitivity ($P_{e}<0.05$, $P_{m}>0.75$, $P_{w}>0.80$ fixed)
   $|\varepsilon|<0.005$ 4.5% 0.05% 99.95% 95.5%
   $|\varepsilon|<0.01$ (selected) 1.9% 0.15% 99.85% 98.1%
   $|\varepsilon|<0.02$ 1.2% 0.32% 99.68% 98.8%
   $|\varepsilon|<0.05$ 1.1% 0.55% 99.45% 98.9%
(c) $P_{m}$ Sensitivity ($P_{e}<0.05$, $|\varepsilon|<0.01$, $P_{w}>0.80$ fixed)
   $P_{m}>0.50$ 2.6% 0.10% 99.90% 97.4%
   $P_{m}>0.65$ 2.1% 0.12% 99.88% 97.9%
   $P_{m}>0.75$ (selected) 1.9% 0.15% 99.85% 98.1%
   $P_{m}>0.85$ 1.6% 0.21% 99.79% 98.4%
   $P_{m}>0.90$ 1.4% 0.35% 99.65% 98.6%
(d) $P_{w}$ Sensitivity ($P_{e}<0.05$, $|\varepsilon|<0.01$, $P_{m}>0.75$ fixed)
   $P_{w}>0.50$ 3.0% 0.07% 99.93% 97.0%
   $P_{w}>0.70$ 2.2% 0.11% 99.89% 97.8%
   $P_{w}>0.80$ (selected) 1.9% 0.15% 99.85% 98.1%
   $P_{w}>0.85$ 1.5% 0.25% 99.75% 98.5%
   $P_{w}>0.90$ 1.3% 0.42% 99.58% 98.7%
Appendix G On-chain Gas Cost Analysis
We implemented the VeriLLM verification smart contract in Solidity and measured gas consumption using Hardhat. Table 8 shows the results on Optimism L2 (gas price: 0.0013 Gwei, ETH: \$3,098 USD).
Table 8. Gas Consumption and Cost on Optimism
Operation Gas Cost (USD)
Create Task 141,466 \$0.00057
Commit Inference 70,532 \$0.00028
Commit Verification 167,100 \$0.00067
Generate Sampling (VRF) 284,740 \$0.00115
Reveal Verification 134,000 \$0.00054
Total 797,838 \$0.0032
Appendix H Open Science
To support the reproducibility of our research, we provide a complete artifact package at an anonymous repository  [^1] Available at: https://anonymous.4open.science/r/verillm-experiments-26E5. This repository includes the full implementation of the VeriLLM protocol, all necessary configuration files, and execution scripts, covering homogeneous and heterogeneous hardware benchmarks as well as various quantization-based security evaluations. Our pre-trained models are hosted on the HuggingFace Hub. For full reproduction, we recommend a system equipped with an NVIDIA GPU (32GB+ VRAM) and an Apple Silicon device to support the cross-vendor and numerical stability tests. Detailed environment setup instructions and expected result benchmarks are documented in the README.md file.